# LQA: A Lightweight Quantized-Adaptive Framework for Vision-Language Models on the Edge

### v1

@@ 1 Introduction
Model quantization offers an orthogonal lever for on-device model inference: representing weights with few-bit integers shrinks the model and accelerates inference (Liu et al., 2025b; Badri & Shaji, 2023). While quantization has been extensively studied for general model inference (Jacob et al., 2018; Gholami et al., 2021; Liu et al., 2025a), its integration with TTA remains largely unexplored due to unique challenges. The effectiveness of most TTA methods relies heavily on backpropagation gradients, but low-precision arithmetic in quantized models increases both the computational cost and instability of backpropagation, often resulting in significant performance degradation (Niu et al., 2024). For backpropagation-free TTA methods, such as the gradient-free adapters in TDA (Karmanov et al., 2024), it is still an open question how reliably the adapter parameters can be optimized when using these algorithms on a quantized backbone. Currently, there is no end-to-end TTA solution that is both robust to distribution shifts and suitable for deployment on resource-constrained devices.
@@ 2.2 Quantization of Vision–Language Models
VLMs are often large and computationally intensive, making deployment on resource-limited devices challenging. Post-training quantization compresses neural networks by representing weights, and sometimes activations, with low-precision numbers, thereby reducing memory footprint and accelerating inference. Generic quantization techniques such as bitsandbytes 8-bit quantization (Dettmers et al., 2022b), GPTQ (Frantar et al., 2023a), and AWQ (Lin et al., 2024), have proven effective for LLMs. However, these methods usually require a calibration dataset, which is often unavailable for edge deployment of VLMs. Data-free approaches, such as HQQ (Badri & Shaji, 2023) and Quanto (Corvoysier et al., 2024), provide more flexible fixed-point formats; however, they have not yet been explored in the context of TTA for VLMs. In contrast, our method is both adaptive and quantization-aware, while also being fully compatible with VLMs.
@@ 3.3.2 Selective Precision Retention
3.4 Quantized Test-Time Adaptation (Q-TTA)
Following TDA conventions, adaptation is performed as:
@@ 4.2.2 Quantization baselines
  • Half-Quadratic Quantization (HQQ) (Badri & Shaji, 2023): A data-free method that optimizes for quantization error using a sparsity-promoting loss. We evaluated HQQ across bit sizes (1, 2, 3, 4, 8) and group sizes (8 to 512).
@@ 5 Results
5.1 Comparison with TTA Methods
(a) Memory vs Accuracy
(b) Latency vs Accuracy
Figure 2: (a) Memory consumption and (b) latencies versus accuracy.
@@ 5.5 Ablation Study
When comparing the quantization methods, the LQA-Lite (4-bit) model demonstrates considerable advantages in efficiency. For instance, the LQA-Lite configuration using both encoders reduces memory consumption by 47.45 MB, a 24.2% decrease, and lowers latency by 3.3 ms relative to its 8-bit counterpart. This significant reduction in resource requirements is achieved with a modest decrease in accuracy of only 1.58%, presenting a favorable trade-off for deployment in environments with limited computational resources. A more detailed ablation study of the 1-bit to 8-bit versions of various methods is presented in the Appendix.
@@ Appendix A Datasets
  • CIFAR10 (Krizhevsky, 2009): A well-known dataset used in vision language benchmark (Hendrycks & Dietterich, 2019; Maharana et al., 2024b; Osowiechi et al., 2024; Maharana et al., 2024a), that consist of 60,000 $32\times 32$ color images in 10 classes, with evenly 6,000 images per class. Each class it-self are completely mutually exclusive to each other. By default, the 6,000 of each class is randomly split into 5,000 training data and 1,000 test data.
@@ B.3 Q-TTA Configuration
In the setting of Q-TTA, the mechanism employs both positive and negative caches to adapt the model during test-time without gradient computations. Due to that, it is essential to carefully tune cache-related hyperparameter, including shot capacity (the number of samples kept in cache), $\alpha$ (influence of positive cache), $\beta$ (influence of negative cache) and filtering thresholds (e.g., entropy and mask thresholds), to ensure effective pseudo label reliability. As these parameters directly influence the quality of the adaptation process, careful finetuning is required to practice. Take into this consideration, we utilized the existing configuration for each dataset from official TDA implementation (Karmanov & co‑authors, 2024), as our Q-TTA build upon TDA and they both share the similar core architecture. These configuration has been proved effective to be effective in prior experiments (Karmanov et al., 2024).
@@ Appendix D Quantization Trade-offs Analysis
Appendix E Ablation Study
In addition to trade-offs analysis, we also conduct an extensive ablation study to evaluate the performance of various quantization methods at different precision levels (4-bit, 8-bit, and low-bits) across three encoder configurations: vision encoder only, text encoder only, and both encoders. The results are summarized in Tables 8 and 9, reporting accuracy, runtime latency, and memory usage for each setting.
Method Vision Text Accuracy (%) Latency (sec) Memory (MB)
BnB (4b) 61.45 0.0216 198.69
HQQ (4b) 64.13 0.0092 184.75
Quanto (4b) 61.12 0.0345 298.40
LQA-Lite(Ours) 64.13 0.0089 184.64
BnB (8b) 65.51 0.0561 252.64
HQQ (8b) 65.31 0.0119 232.19
Quanto (8b) 65.80 0.0215 296.65
LQA(Ours) 65.89 0.0113 231.08
BnB (4b) 64.26 0.0107 294.94
HQQ (4b) 64.64 0.0071 248.76
Quanto (4b) 65.42 0.0069 294.40
LQA-Lite(Ours) 65.69 0.0071 265.80
BnB (8b) 65.46 0.0096 278.49
HQQ (8b) 65.52 0.0070 266.75
Quanto (8b) 65.76 0.0068 296.54
LQA(Ours) 65.76 0.0068 296.38
BnB (4b) 59.99 0.0223 189.24
HQQ (4b) 62.91 0.0086 132.67
Quanto (4b) 60.42 0.0372 290.13
LQA-Lite(Ours) 63.98 0.0081 148.31
BnB (8b) 65.50 0.0573 222.34
HQQ (8b) 65.38 0.0116 202.24
Quanto (8b) 65.69 0.0212 291.52
LQA(Ours) 65.56 0.0114 195.76
Method Vision Text Accuracy (%) Latency (sec) Memory (MB)
HQQ (1b) 2.87 0.0273 199.71
HQQ (2b) 46.89 0.0229 211.83
Quanto (2b) 10.09 0.0319 237.72
HQQ (3b) 61.85 0.0343 215.25
HQQ (1b) 3.13 0.0089 257.86
HQQ (2b) 61.28 0.0077 262.60
Quanto (2b) 32.68 0.0060 239.53
HQQ (3b) 63.91 0.0093 262.60
HQQ (1b) 3.02 0.0285 149.46
HQQ (2b) 42.54 0.0195 164.35
Quanto (2b) 9.50 0.0144 102.22
HQQ (3b) 60.59 0.0321 170.66

### v3

@@ 1 Introduction
Model quantization offers an orthogonal lever for on-device model inference: representing weights with few-bit integers shrinks the model and accelerates inference (Liu et al., 2025b; Badri and Shaji, 2023). While quantization has been extensively studied for general model inference (Jacob et al., 2018; Gholami et al., 2021; Liu et al., 2025a), its integration with TTA remains largely unexplored due to unique challenges. The effectiveness of most TTA methods relies heavily on backpropagation gradients, but low-precision arithmetic in quantized models increases both the computational cost and instability of backpropagation, often resulting in significant performance degradation (Niu et al., 2024). For backpropagation-free TTA methods, such as the gradient-free adapters in TDA (Karmanov et al., 2024), it is still an open question how reliably the adapter parameters can be optimized when using these algorithms on a quantized backbone. Currently, there is no end-to-end TTA solution that is both robust to distribution shifts and suitable for deployment on resource-constrained devices.
@@ 2.2 Quantization of Vision–Language Models
VLMs are often large and computationally intensive, making deployment on resource-limited devices challenging. Post-training quantization compresses neural networks by representing weights, and sometimes activations, with low-precision numbers, thereby reducing memory footprint and accelerating inference. Generic quantization techniques such as bitsandbytes 8-bit quantization (Dettmers et al., 2022b), GPTQ (Frantar et al., 2023a), and AWQ (Lin et al., 2024), have proven effective for LLMs. However, these methods usually require a calibration dataset, which is often unavailable for edge deployment of VLMs. Data-free approaches, such as HQQ (Badri and Shaji, 2023) and Quanto (Corvoysier et al., 2024), provide more flexible fixed-point formats; however, they have not yet been explored in the context of TTA for VLMs. In contrast, our method is both adaptive and quantization-aware, while also being fully compatible with VLMs.
@@ 3.3.2 Selective Precision Retention
3.4 Quantized Test-Time Adaptation (Q-TTA)
Following TDA conventions, adaptation is performed as:
@@ 4.2.2 Quantization baselines
  • Half-Quadratic Quantization (HQQ) (Badri and Shaji, 2023): A data-free method that optimizes for quantization error using a sparsity-promoting loss. We evaluated HQQ across bit sizes (1, 2, 3, 4, 8) and group sizes (8 to 512).
@@ 5.1 Comparison with TTA Methods
(a) Memory vs Accuracy
(b) Latency vs Accuracy
Figure 3: (a) Memory consumption and (b) latencies versus accuracy.
@@ 5.5 Ablation Study
When comparing the quantization methods, the LQA-Lite (4-bit) model demonstrates considerable advantages in efficiency. For instance, the LQA-Lite configuration using both encoders reduces memory consumption by 47.45 MB, a 24.2% decrease, and lowers latency by 3.3 ms relative to its 8-bit counterpart. This significant reduction in resource requirements is achieved with a modest decrease in accuracy of only 1.58%, presenting a favorable trade-off for deployment in environments with limited computational resources.
In addition to trade-offs analysis, we also conduct an extensive ablation study to evaluate the performance of various quantization methods at different precision levels (4-bit, 8-bit, and low-bits) across three encoder configurations: vision encoder only, text encoder only, and both encoders. The results are summarized in Tables 6 and 7, reporting accuracy, runtime latency, and memory usage for each setting.
Method Vision Text Accuracy (%) Latency (sec) Memory (MB)
BnB (4b) 61.45 0.0216 198.69
HQQ (4b) 64.13 0.0092 184.75
Quanto (4b) 61.12 0.0345 298.40
LQA-Lite(Ours) 64.13 0.0089 184.64
BnB (8b) 65.51 0.0561 252.64
HQQ (8b) 65.31 0.0119 232.19
Quanto (8b) 65.80 0.0215 296.65
LQA(Ours) 65.89 0.0113 231.08
BnB (4b) 64.26 0.0107 294.94
HQQ (4b) 64.64 0.0071 248.76
Quanto (4b) 65.42 0.0069 294.40
LQA-Lite(Ours) 65.69 0.0071 265.80
BnB (8b) 65.46 0.0096 278.49
HQQ (8b) 65.52 0.0070 266.75
Quanto (8b) 65.76 0.0068 296.54
LQA(Ours) 65.76 0.0068 296.38
BnB (4b) 59.99 0.0223 189.24
HQQ (4b) 62.91 0.0086 132.67
Quanto (4b) 60.42 0.0372 290.13
LQA-Lite(Ours) 63.98 0.0081 148.31
BnB (8b) 65.50 0.0573 222.34
HQQ (8b) 65.38 0.0116 202.24
Quanto (8b) 65.69 0.0212 291.52
LQA(Ours) 65.56 0.0114 195.76
Method Vision Text Accuracy (%) Latency (sec) Memory (MB)
HQQ (1b) 2.87 0.0273 199.71
HQQ (2b) 46.89 0.0229 211.83
Quanto (2b) 10.09 0.0319 237.72
HQQ (3b) 61.85 0.0343 215.25
HQQ (1b) 3.13 0.0089 257.86
HQQ (2b) 61.28 0.0077 262.60
Quanto (2b) 32.68 0.0060 239.53
HQQ (3b) 63.91 0.0093 262.60
HQQ (1b) 3.02 0.0285 149.46
HQQ (2b) 42.54 0.0195 164.35
Quanto (2b) 9.50 0.0144 102.22
HQQ (3b) 60.59 0.0321 170.66
@@ Appendix A Datasets
  • CIFAR10 (Krizhevsky, 2009): A well-known dataset used in vision language benchmark (Hendrycks and Dietterich, 2019; Maharana et al., 2024a; Osowiechi et al., 2024; Maharana et al., 2024b), that consist of 60,000 $32\times 32$ color images in 10 classes, with evenly 6,000 images per class. Each class it-self are completely mutually exclusive to each other. By default, the 6,000 of each class is randomly split into 5,000 training data and 1,000 test data.
@@ B.3 Q-TTA Configuration
In the setting of Q-TTA, the mechanism employs both positive and negative caches to adapt the model during test-time without gradient computations. Due to that, it is essential to carefully tune cache-related hyperparameter, including shot capacity (the number of samples kept in cache), $\alpha$ (influence of positive cache), $\beta$ (influence of negative cache) and filtering thresholds (e.g., entropy and mask thresholds), to ensure effective pseudo label reliability. As these parameters directly influence the quality of the adaptation process, careful finetuning is required to practice. Take into this consideration, we utilized the existing configuration for each dataset from official TDA implementation (Karmanov and co‑authors, 2024), as our Q-TTA build upon TDA and they both share the similar core architecture. These configuration has been proved effective to be effective in prior experiments (Karmanov et al., 2024).